# Robust Barycenters of Persistence Diagrams

### v1

@@ II-B Wasserstein distance and barycenter
where $\phi:X\to Y$ is a bijection between $X$ and $Y$. The transportation cost $c_{q}$, based on powered distances, is such that $c_{q}(x,y)=\|x-y\|_{2}^{q}$ if $x\notin\Delta$ or $y\notin\Delta$, and 0 otherwise. An optimal bijection $\phi^{*}$ minimizing Eq. 1 is called an optimal transport plan. In practice, $q$ is often set to $2$, yielding the Wasserstein distance, noted $W_{2}$. A Wasserstein $W_{q}$ barycenter of persistence diagrams $X_{1},\ldots,X_{m}$[78], is defined by minimizing the Fréchet energy:
$\operatorname*{arg\,min}_{B}\displaystyle\sum\limits_{i=1}^{m}\lambda_{i}W_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}q}}^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}q}}(B,X_{i}),$ (2)
where $\lambda_{i}\geq 0$ and $\sum_{i}\lambda_{i}=1$. Practical algorithms have been proposed for the computation of Wasserstein barycenters [78, 80], but only for the specific case where $q=2$. When $q=2$, a solution $B^{*}$ of Eq. 2 has the following property: a point $x\in B^{*}$ is an arithmetic mean of $m$ points each in $X_{1},\ldots,X_{m}$, which considerably eases the optimization of Eq. 2. Fig. 2 (center) presents an example of a $W_{2}$ barycenter. A barycenter computed using the $W_{2}$ distance emulates the behavior of a mean of an ensemble of scalars. As such, it is prone to the influence of outliers. This motivates the use of a more general framework for computing robust barycenters (detailed in Sec. III).
@@
    Compute $\phi_{i}\in\displaystyle\operatorname*{arg\,min}_{\phi:B^{(t)}\to X_{i}}\sum\limits_{\ell=1}^{K}c_{q}\big(x_{\ell}^{(t)},\phi(x_{\ell}^{(t)})\big)$.
@@
     Find $x_{\ell}^{(t+1)}=\mathfrak{b}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}q}}\big(\phi_{1}(x_{\ell}^{(t)}),\ldots,\phi_{m}(x_{\ell}^{(t)})\big)$.
@@ III-A Optimization
$\mathfrak{b}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}q}}:\begin{cases}\mathbb{R}^{2}\times\ldots\times\mathbb{R}^{2}\to\mathbb{R}^{2}\\ (y_{1},\ldots,y_{m})\mapsto\displaystyle\operatorname*{arg\,min}_{x}\sum\limits_{i=1}^{m}\lambda_{i}c_{q}(x,y_{i})\end{cases}.$ (3)
To optimize Eq. 3, we use gradient descent, by leveraging the automatic differentiation capabilities of PyTorch. This optimization procedure is plugged into Alg. 1 in the update step (ground barycenter computation line). In practice we give a maximum number $T$ of overall iterations. We noticed that taking $T<10$ is sufficient for convergence. Assumptions for achieving convergence are discussed in the next section.
Figure 3: Simple example where computing the arithmetic mean instead of optimizing $\mathfrak{b}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}q}}$ increases the Fréchet energy (noted $E_{F}$) for $q=1$. We have three simple persistence diagrams, in dark blue, gray and yellow, each having a single point. For this problem, the transport plans are fixed and the barycenter has only one point. On the left we initialized the barycenter as the diagram encoded in green. In the middle, we have the candidate of the barycenter encoded in cyan when computing an arithmetic mean after one iteration. We can see that the Fréchet energy (for $q=1$) increased. On the right, we have a candidate for the barycenter encoded in purple when optimizing $\mathfrak{b}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}q}}$ instead, this time displaying a decrease of the Fréchet energy at one iteration.
@@ III-B Convergence
Figure 4: Comparison of clustering results on an ensemble of diagrams of Gaussian mixtures. On the left we have the 3 clusters: one cluster of 2 Gaussians (top), one cluster of 3 Gaussians (middle) and one cluster of 4 Gaussians (bottom). In the first and second clusters, we inserted an outlier (highlighted in green and cyan respectively) by setting one isolated pixel to an arbitrarily high value. Those pixels result in persistent pairs in the corresponding diagrams. On top we have the distance matrices of $W_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}q}}$ for $q\in\{2,1.8,1.6,1.4,1.2,1\}$. In the distance matrices, the clustering results are shown with dashed squares (clusters are colored in dark purple, purple and pale purple) while the outlier diagrams are indicated with a plain square (green and cyan). In the three frames; we visualize the evolution of each cluster and their barycenters for each $q$. Each frame corresponds to a cluster (top: cluster 1, middle: cluster 2, bottom: cluster 3). The outlier diagrams are colored in green and cyan. The barycenters are shown in opaque while the diagrams of each cluster are shown in transparent. We observe that for $q\in\{2,1.8\}$ the green outlier is incorrectly assigned to the second cluster (as it exhibits the same number of persistence pairs, 3, as the entries of cluster 2). Similarly, given its number of persistence pairs, the cyan outlier is incorrectly assigned to the third cluster until $q=1.6$. Beyond this value, the effect of the extra feature in these outlier diagrams is decreased, enabling their correct clustering.
Tanguy et al. prove, in the setup of probability measures, that a fixed-point method for minimizing Eq. 3 converged under certain assumptions [74]. In this section, we review these assumptions in the setup of persistence diagrams to argue the convergence of our overall approach.
@@
For all $(y_{1},\ldots,y_{m})\in\mathbb{R}^{2}\times\ldots\times\mathbb{R}^{2}$, for all $\lambda_{1},\ldots,\lambda_{m}$ barycentric coefficients, $\operatorname*{arg\,min}_{x}\sum\limits_{i=1}^{m}\lambda_{i}c_{q}(x,y_{i})$ is reduced to a single element.
@@
We prove that this assumption is satisfied in our case in Appendix A, when $q>1$. When $q=1$, this assumption does not hold in general (due to the presence of collinear points, in particular on the diagonal). In practice, this may lead to numerical instabilities when considering $q=1$, especially when ground barycenters are not unique.
Under those assumptions, and denoting the Fréchet energy $E_{F}(B)=\sum\limits_{i=1}^{m}\lambda_{i}W^{q}_{q}(B,X_{i})$, Tanguy et al. [74] show that for two consecutive iterates $B^{(t)}$ and $B^{(t+1)}$, we have $E_{F}(B^{(t)})\geq E_{F}(B^{(t+1)})$. This means that fixed-point iterations $(B^{(t)})$ decrease the energy optimized in Eq. 2.
For $q\in(1,2)$, a resulting fixed point is a barycenter that is more robust to the presence of an outlier, in the initial set $X_{1},\ldots,X_{m}$, compared to a $W_{2}$ barycenter. Fig. 2 illustrates this difference when computing a barycenter with an outlier.
@@
Figure 5: Visual comparison of distance matrices using $W_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}q}}$ for $q\in\{2,1.8,1.6,1.4,1.2,1\}$ on the Volcanic Eruption ensemble and the clustering results. Distance matrices are represented similarly to Fig. 4. This ensemble of 12 persistence diagrams has a natural outlier highlighted in cyan on the distance matrices. On the top, we can see that for $q\in\{2,1.8\}$, the clustering algorithm keeps the outlier alone, groups the 8 first diagrams together and groups the last three together. Then starting from 1.6 to 1.2, the correct clusters are returned. But for $q=1$, we can see that the clusters are not discriminated enough. On the bottom we have one representative scalar field for each cluster, and on the bottom left the corresponding diagrams, the cyan scalar field and diagram being the outlier. On the right, we have a visual comparison of two barycenters of the last cluster of four diagrams (represented on the bottom right of the square). The pink one encodes a $W_{2}$ barycenter, while the green one encodes a $W_{1.2}$ barycenter. We can see the influence of the outlier in the pink barycenter, as the two global pairs are higher than the green ones testifying the difference of scaling between the outlier (cyan) and the three other diagrams in the cluster. Also, we notice the presence of an isolated pair above the diagonal (pink diagram, rightmost pair) that is generated by the isolated persistent pair in the outlier diagram (cyan diagram, rightmost pair). This outlier pair is not present in the $W_{1.2}$ barycenter (green diagram), hence making it more visually representative of the ensemble than the $W_{2}$ barycenter (pink).
@@ IV Results
This section presents two applications of our robust barycenters (Sec. III) along with detailed experiments. The experimental results are obtained on a computer with an NVIDIA Geforce RTX 2060 (Mobile Q) with 6 GB of dedicated VRAM. Our methods were implemented on Python, using Pytorch for computations on the GPU. We ran some experiments on two public ensembles [62] described in [63]. One is an acquired 2D ensemble and the other is a simulated 3D ensemble, both selected from past SciVis contests [59]. For the experiments, only the persistence pairs containing maxima were considered.
@@ IV-A Clustering on the persistence diagram metric space
For this, we consider the classic clustering method, the $k$-means algorithm. This is an iterative algorithm alternating between two phases: computing k barycenters and labeling the elements into k clusters. At first, $k$ cluster barycenters $B_{j}$ with $j\in\{1,\ldots,k\}$ are initialized as $k$ diagrams in the initial input set $X_{1},\ldots,X_{N}$, typically using the $k$-means++ method [17]. Then, the labeling phase consists in assigning each diagrams $X_{n}$ to the closest barycenter $B_{j}$ by using the $W_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}q}}$ distance. After the labeling phase, the barycenters are updated by computing new barycenters based on the new $k$ clusters using Alg. 1. The algorithm stops when reaching a maximum number of iterations or when converging (i.e., when the labels do not change anymore). However, when using $W_{2}$ as a distance, the presence of an outlier in an ensemble of persistence diagrams can incorrectly influence the barycenters in the $k$-means algorithm, and as a consequence the output labels of the clustering.
@@
Figure 6: Visual comparison of 2D planar layouts (on the right) of Wasserstein barycenters after solving Eq. 4 and Eq. 5 when taking a dictionary with 3 diagrams. The initial ensemble Isabel is composed of 12 diagrams divided in 3 classes of 4 diagrams each. We removed 2 diagrams from the first and second classes, yielding an imbalanced ensemble (in terms of class size) of 8 diagrams. We show representative scalar field for each cluster, along with their diagram. The clusters are colored in dark purple, purple and cyan respectively. In the planar layouts, the points, representing the barycenters, are colored by their ground truth classification. For $q=2$, the barycenter approximating an element of the first cluster (dark purple) is misplaced (i.e., located near the second cluster, purple). For $q=1.5$, this same barycenter is correctly placed, thus yielding a planar projection that is more faithful to the ground-truth classification.
@@ IV-B Wasserstein dictionary encoding of persistence diagrams
Another application consists in using our robust barycenters as a core procedure for dictionary based encodings of ensembles of persistence diagrams [71]. Let $X_{1},\ldots,X_{N}$ be an ensemble of persistence diagrams. A Wasserstein dictionary encoding aims at optimizing a set of persistence diagrams $\mathcal{D}^{*}=\{a_{1}^{*},\ldots,a_{m}^{*}\}$ (called dictionary) and $N$ vectors of barycentric coefficients $\Lambda^{*}=\{\bm{\lambda}_{1}^{*},\ldots,\bm{\lambda}_{N}^{*}\}$ (i.e., $N$ vectors of size $m$, with positive elements summing to 1) by solving:
$\operatorname*{arg\,min}_{\mathcal{D},\Lambda}\displaystyle\sum\limits_{\ell=1}^{N}W_{2}^{2}\big(B_{2}(\mathcal{D},\bm{\lambda}_{\ell}),X_{\ell}\big),$ (4)
@@
$\operatorname*{arg\,min}_{\mathcal{D},\Lambda}\displaystyle\sum\limits_{l=1}^{N}W_{q}^{q}\big(B_{q}(\mathcal{D},\bm{\lambda}_{l}),X_{l}\big),$ (5)

### v2

@@ II-B Wasserstein distance and barycenter
where $\phi:X\to Y$ is a bijection between $X$ and $Y$. The transportation cost $c_{q}$, based on powered distances, is such that $c_{q}(x,y)=\|x-y\|_{2}^{q}$ if $x\notin\Delta$ or $y\notin\Delta$, and 0 otherwise. An optimal bijection $\phi^{*}$ minimizing Eq. 1 is called an optimal transport plan. In practice, $q$ is often set to $2$, yielding the 2-Wasserstein distance, noted $W_{2}$. For $q\geq 1$, the $q$-Wasserstein distance defines a metric on the space of persistence diagrams [79]. A Wasserstein $W_{q}$ barycenter of persistence diagrams $X_{1},\ldots,X_{m}$[80], is defined by minimizing the Fréchet energy over the space of persistence diagrams $\mathbb{D}$:
$\operatorname*{arg\,min}_{{B\in\mathbb{D}}}\displaystyle\sum\limits_{i=1}^{m}\lambda_{i}W_{{q}}^{{q}}(B,X_{i}),$ (2)
where $\lambda_{i}\geq 0$ and $\sum_{i}\lambda_{i}=1$ (i.e $\bm{\lambda}=(\lambda_{1},\ldots,\lambda_{m})\in\Sigma_{m}$). Practical algorithms have been proposed for the computation of Wasserstein barycenters [80, 81], but only for the specific case where $q=2$. When $q=2$, a solution $B^{*}$ of Eq. 2 has the following property: a point $x\in B^{*}$ is an arithmetic mean of $m$ points each in $X_{1},\ldots,X_{m}$, which considerably eases the optimization of Eq. 2. Fig. 2 (center) presents an example of a $W_{2}$ barycenter. A barycenter computed using the $W_{2}$ distance emulates the behavior of a mean of an ensemble of scalars. As such, it is prone to the influence of outliers. This motivates the use of a more general framework for computing robust barycenters (detailed in Sec. III).
@@
      Compute $\phi_{i}\leftarrow\displaystyle\operatorname*{arg\,min}_{\phi:B^{(t)}\to X_{i}}\sum\limits_{\ell=1}^{K}c_{q}\big(x_{\ell}^{(t)},\phi(x_{\ell}^{(t)})\big)$.
@@
     Find $x_{\ell}^{(t+1)}=\mathfrak{b}_{{q}}\big(\phi_{1}(x_{\ell}^{(t)}),\ldots,\phi_{m}(x_{\ell}^{(t)})\big)$.
@@ III-A Optimization
$\mathfrak{b}_{{q}}:\begin{cases}\mathbb{R}^{2}\times\ldots\times\mathbb{R}^{2}\to\mathbb{R}^{2}\\ (y_{1},\ldots,y_{m})\mapsto\displaystyle\operatorname*{arg\,min}_{{x\in\mathbb{R}^{2}}}\sum\limits_{i=1}^{m}\lambda_{i}c_{q}(x,y_{i})\end{cases}.$ (3)
To optimize Eq. 3, we use gradient descent, by leveraging the automatic differentiation capabilities of PyTorch (for which we give the parameters used in Appendix A). This optimization procedure is plugged into Alg. 1 in the update step (ground barycenter computation line). In practice we give a maximum number $T$ of overall iterations. Regarding the time complexity of Alg. 1, computing an optimal matching between two persistence diagrams using traditional linear programming methods requires a runtime complexity of $\mathcal{O}(K^{3})$ [63]. Also, to update the barycenter, a gradient descent scheme to compute $\mathfrak{b}_{q}$ has a complexity of $\mathcal{O}(mT^{\prime})$[57] with $T^{\prime}\leq 250$ the maximum number of iterations for the optimization of Eq. 3. Thus, overall Alg. 1 has a runtime complexity of $\mathcal{O}\left(T(mK^{3}+mT^{\prime}K)\right)$ in practice. We noticed that taking $T<10$ is sufficient for convergence. Assumptions for achieving convergence are discussed in the next section.
Figure 3: Simple example where computing the arithmetic mean instead of optimizing $\mathfrak{b}_{{q}}$ increases the Fréchet energy (noted $E_{F}$) for $q=1$. We have three simple persistence diagrams, in dark blue, gray and yellow, each having a single point. For this problem, the transport plans are fixed and the barycenter has only one point. On the left we initialized the barycenter as the diagram encoded in green. In the middle, we have the candidate of the barycenter encoded in cyan when computing an arithmetic mean after one iteration. We can see that the Fréchet energy (for $q=1$) increased. On the right, we have a candidate for the barycenter encoded in purple when optimizing $\mathfrak{b}_{{q}}$ instead, this time displaying a decrease of the Fréchet energy at one iteration.
@@ III-B Convergence
Figure 4: Comparison of clustering results on an ensemble of diagrams of Gaussian mixtures. On the left we have the 3 clusters: one cluster of 2 Gaussians (top), one cluster of 3 Gaussians (middle) and one cluster of 4 Gaussians (bottom). In the first and second clusters, we inserted an outlier (highlighted in green and cyan respectively) by setting one isolated pixel (highlighted in red) to an arbitrarily high value. Those pixels result in persistent pairs in the corresponding diagrams. On top we have the distance matrices of $W_{{q}}$ for $q\in\{2,1.8,1.6,1.4,1.2,1\}$. In the distance matrices, the clustering results are shown with dashed squares (clusters are colored in dark purple, purple and pale purple) while the outlier diagrams are indicated with a plain square (green and cyan). In the three frames; we visualize the evolution of each cluster and their barycenters for each $q$. Each frame corresponds to a cluster (top: cluster 1, middle: cluster 2, bottom: cluster 3). The outlier diagrams are colored in green and cyan. The barycenters are shown in opaque while the diagrams of each cluster are shown in transparent. We observe that for $q\in\{2,1.8\}$ the green outlier is incorrectly assigned to the second cluster (as it exhibits the same number of persistence pairs, 3, as the entries of cluster 2), then, from $q=1.6$, it is correctly assigned to the first cluster as shown by the green arrow (it shifts from the second frame to the first one). Similarly, given its number of persistence pairs, the cyan outlier is incorrectly assigned to the third cluster until $q=1.6$, then from $q=1.4$ it shifts to the second cluster as indicated by the cyan arrow.
Tanguy et al. [76] prove, in the setup of probability measures, that a fixed-point method for minimizing Eq. 3 converged under certain assumptions [76]. In this section, we review these assumptions in the setup of persistence diagrams to argue the convergence of our overall approach.
@@
For all $(y_{1},\ldots,y_{m})\in\mathbb{R}^{2}\times\ldots\times\mathbb{R}^{2}$, for all $\lambda_{1},\ldots,\lambda_{m}$ barycentric coefficients, $\operatorname*{arg\,min}_{{x\in\mathbb{R}^{2}}}\sum\limits_{i=1}^{m}\lambda_{i}c_{q}(x,y_{i})$ is reduced to a single element.
@@
We prove that this assumption is satisfied in our case in Appendix B, when $q>1$. When $q=1$, this assumption does not hold in general (due to the presence of collinear points, in particular on the diagonal). In practice, this may lead to numerical instabilities when considering $q=1$, especially when ground barycenters are not unique. Thus, our approach does not support the case $q=1$ in general and we recommend to take $q>1$ (unless the input diagrams are simple and not prone to numerical instabilities, e.g. Fig. 4).
Under those assumptions, and denoting the Fréchet energy $E_{F}(B)=\sum\limits_{i=1}^{m}\lambda_{i}W^{q}_{q}(B,X_{i})$, Tanguy et al. [76] show that for two consecutive iterates $B^{(t)}$ and $B^{(t+1)}$, we have $E_{F}(B^{(t)})\geq E_{F}(B^{(t+1)})$. This means that fixed-point iterations $(B^{(t)})$ decrease the energy optimized in Eq. 2. However, while the iterates decrease $E_{F}$, it is not guaranteed that the obtained fixed-point is a global (even local) minimum of $E_{F}$. For $q\in(1,2)$, a resulting fixed point is a barycenter that is more robust to the presence of an outlier, in the initial set $X_{1},\ldots,X_{m}$, compared to a $W_{2}$ barycenter. Fig. 2 illustrates this difference when computing a barycenter with an outlier. Basically, the closer $q$ is to 2, the more it behaves like a mean. Conversely, the closer $q$ is to 1, the more it behaves like a median (i.e., describing typical central values, less sensitive to outliers than averages).
@@
Figure 5: Visual comparison of distance matrices using $W_{{q}}$ for $q\in\{2,1.8,1.6,1.4,1.2,1\}$ on the Volcanic Eruption ensemble and the clustering results. This ensemble of 12 persistence diagrams has a natural outlier highlighted in cyan on the distance matrices. On the top, we can see that for $q\in\{2,1.8\}$, the clustering algorithm keeps the outlier alone, groups the 8 first diagrams together and groups the last three together. Then starting from 1.6 to 1.2, the correct clusters are returned. But for $q=1$, we can see that the clusters are not discriminated enough. In the middle we have one representative scalar field for each cluster, and on the right the corresponding diagrams, the cyan scalar field and diagram being the outlier. On the bottom, we have a visual comparison of three barycenters (pink) of the last cluster in three cases: one computed with the outlier in cyan when $q=2$, one without the outlier when $q=2$ and one with the outlier when $q=1.2$. We witness the influence of the outlier on the barycenter on the left, as there are two persistent pairs (in pink) higher than the ones in the other three diagrams of the cluster. Also, we notice the presence of an isolated pair above the diagonal that is generated by the isolated cyan pair. In the other cases, the barycenters (pink) are very similar, showing the robustness of this $1.2$-barycenter to the presence of this outlier.
@@ IV Results
This section presents two applications of our robust barycenters (Sec. III) along with detailed experiments. The experimental results are obtained on a computer with an NVIDIA Geforce RTX 2060 (Mobile Q) with 6 GB of dedicated VRAM. Our methods were implemented on Python, using Pytorch for computations on the GPU. Given the specification of our GPU, we were memory-limited in our experiments, given the modest dedicated VRAM. This led us to run experiments with diagrams thresholded by persistence, to reach a typical size of $\sim$20 points per diagram (as a reminder, in practice, all diagrams need to be augmented, Sec. II-A). We ran some experiments on two public ensembles [64] described in [65]. One is an acquired 2D ensemble and the other is a simulated 3D ensemble, both selected from past SciVis contests [60]. For the experiments, only the persistence pairs containing maxima were considered.
@@ IV-A Clustering on the persistence diagram metric space
For this, we consider the classic clustering method, the $k$-means algorithm. This is an iterative algorithm alternating between two phases: computing k barycenters and labeling the elements into k clusters. At first, $k$ cluster barycenters $B_{j}$ with $j\in\{1,\ldots,k\}$ are initialized as $k$ diagrams in the initial input set $X_{1},\ldots,X_{N}$, typically using the $k$-means++ method [17]. Then, the labeling phase consists in assigning each diagrams $X_{n}$ to the closest barycenter $B_{j}$ by using the $W_{{q}}$ distance. After the labeling phase, the barycenters are updated by computing new barycenters based on the new $k$ clusters using Alg. 1. The algorithm stops when reaching a maximum number of iterations or when converging (i.e., when the labels do not change anymore). However, when using $W_{2}$ as a distance, the presence of an outlier in an ensemble of persistence diagrams can incorrectly influence the barycenters in the $k$-means algorithm, and as a consequence the output labels of the clustering.
@@
Figure 6: Visual comparison of 2D planar layouts of Wasserstein barycenters generated by the Wasserstein dictionary encoding method with $q=2$ (top-left) and $q=1.5$ (bottom-left). The initial Isabel ensemble is divided in 3 classes of 4 diagrams, and we removed 4 entries to create an imbalance in cluster sizes. We show a representative scalar field and diagram for each cluster highlighted in three different colors. In the planar layouts, the points, representing the barycenters, are colored by their ground truth classification. For $q=2$, the method misplaces a point from the first cluster (dark purple) near the purple points. For $q=1.5$, the method provides well separated classes, thus yielding a planar projection that is more faithful to the ground-truth classification.
@@ IV-B Wasserstein dictionary encoding of persistence diagrams
Another application consists in using our robust barycenters as a core procedure for dictionary based encodings of ensembles of persistence diagrams [73]. Let $X_{1},\ldots,X_{N}$ be an ensemble of persistence diagrams. A Wasserstein dictionary encoding aims at optimizing a set of persistence diagrams $\mathcal{D}^{*}=\{a_{1}^{*},\ldots,a_{m}^{*}\}$ (called dictionary) and $N$ vectors of barycentric coefficients $\Lambda^{*}=\{\bm{\lambda}_{1}^{*},\ldots,\bm{\lambda}_{N}^{*}\}$$\in\Sigma_{m}^{N}$ by solving:
$\operatorname*{arg\,min}_{{\mathcal{D}\in\mathbb{D}^{m},\Lambda\in\Sigma_{m}^{N}}}\displaystyle\sum\limits_{\ell=1}^{N}W_{2}^{2}\big(B_{2}(\mathcal{D},\bm{\lambda}_{\ell}),X_{\ell}\big),$ (4)
@@
$\operatorname*{arg\,min}_{{\mathcal{D}\in\mathbb{D}^{m},\Lambda\in\Sigma_{m}^{N}}}\displaystyle\sum\limits_{l=1}^{N}W_{q}^{q}\big(B_{q}(\mathcal{D},\bm{\lambda}_{l}),X_{l}\big),$ (5)